# Equilibrium Portfolio Selection under Utility-Variance Analysis of Log Returns in Incomplete Markets

### v1

@@ 3 A Necessary Condition of Time-Consistent Equilibria
Consequently, the desired result holds that
$\sigma\hat{\pi}=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}.$
@@ 4 Study of BSDE System in the Incomplete Factor Model
In this section, we focus on the factor model and the asssociated BSDEs (3.1). The factor model provides an incomplete market setting with stochastic parameters in which the risk-free rate, the stock return rate, and the volatility rate can be expressed by a deterministic function of time $t$ and the factor $X_{t}$. It is assumed that the stochastic factor $X$ is governed by
@@
$\mu(t,x),r(t,x),\sigma(t,x)>0$ and $\theta(t,x)=\sigma^{-1}(\mu-r)(t,x)$ are bounded measurable (w.r.t. $\mathcal{B}([0,T]\times\mathbb{R})$) deterministic functions with respect to $(t,x)$. Here, $\theta$ represents the market price of risk, which is assumed to be positive and bounded.
@@ 3. There exists a constant $L$ such that
Under Assumption 4.2, SDE (4.1) admits a unique strong solution. It is noted that the log-return process $R$ does not appear in the BSDE system (3.1), which motivates us to consider the solution adapted to the filtration $\mathbb{F}_{\bar{B}}$. We can therefore conjecture that $Z^{1}=\tilde{Z}^{1}=0$ and aim to find the solution to the following Markovian BSDE system.
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+\nu(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|\rho Z_{s}+\hat{u}_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\hat{\pi}_{s})\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &Y_{T}=0,\quad\tilde{Y}_{T}=0,\\ &\hat{u}=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma},\hat{\pi}=\sigma^{-1}\hat{u},a(\cdot,\pi):=r+(\mu-r)\pi-\frac{1}{2}|\sigma\pi|^{2}.\end{aligned}\right$. (4.2)
@@ 4.1 Existence of time-consistent equilibrium when $\rho=0$
In this subsection, we consider the case $\rho=0$. Then it follows that
$\hat{\pi}_{t}=\sigma_{t}^{-1}\hat{u}_{t}=\sigma_{t}^{-1}\frac{e^{-\zeta Y_{t}}\theta(t,X_{t})}{(\zeta+1)e^{-\zeta Y_{t}}+\gamma},$ (4.3)
and the BSDE system (3.1) can be decoupled as
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+\nu(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y_{s}=\left(\frac{\zeta+1}{2}\hat{u}_{s}^{2}+\frac{\zeta}{2}|Z_{s}|^{2}-r(s,X_{s})-\theta(s,X_{s})\hat{u}_{s}\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &X_{0}=x,\quad Y_{T}=0.\end{aligned}\right$. (4.4)
@@
$\left\{\begin{aligned} &\mathrm{d}\tilde{Y}_{s}=-r(s,X_{s})-\theta(s,X_{s})\hat{u}_{s}+\frac{1}{2}|\hat{u}_{t}|^{2}\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &\tilde{Y}_{T}=0.\end{aligned}\right$. (4.5)
It is straightforward to see that one can first solve equation (4.4) and then substitute this solution into equation (4.5) to obtain the solution for the entire system. Regarding the existence of the solution to BSDE (4.4), we verify conditions (2A3), (3A2), (2A1), and (4A2) and apply Theorem 4.3 in FAN20161511, which leads to the following existence and uniqueness result.
@@
If Assumptions 4.1 and 4.2 hold, the BSDE (4.4) has a unique solution $(Y,Z)\in\mathcal{S}^{\infty}\times L^{2}_{\mathbb{F}}(0,T;\mathbb{R})$.
@@
In the following, we establish that $\hat{\pi}$ is an equilibrium strategy by means of BSDEs (4.5) and (4.3), along with Lemma 3.4. This result is stated in the theorem below.
@@
When $\rho=0$, suppose that Assumptions 4.1 and 4.2 hold, and let $Y$ denote the bounded solution to BSDE (4.4). Then the strategy defined in (4.3) is an equilibrium strategy.
@@
For notational simplicity, let $\hat{R}=R^{\hat{\pi}}$ denote the log-return process associated with $\hat{\pi}$. Moreover, we use $\pi^{t,\varepsilon}$ and $\xi^{t,\varepsilon}$ to represent the perturbation strategy $\hat{\pi}^{t,\varepsilon,\eta}$ and the corresponding difference process for any fixed $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$ and $\varepsilon\in(0,T-t)$.
We first verify that $\hat{\pi}$ is an admissible strategy. Since $\hat{\pi}$ is bounded, it follows that $\hat{\pi}\in H_{\textup{BMO}}$. It is straightforward to show that
$\mathbb{E}\left[\exp\left(c\left|\int_{t}^{T}|\sigma_{s}\hat{\pi}_{s}|\mathrm{d}B_{s}\right|\right)\right]\leq 2\exp\left(\frac{c^{2}}{2}\|\sigma\hat{\pi}\|_{\infty}^{2}(T-t)\right),\quad\forall c>0.$
Combining this with $\hat{R}_{T}-\hat{R}_{t}=\int_{t}^{T}\left(r_{s}+(\mu_{s}-r_{s})\hat{\pi}_{s}-\frac{1}{2}|\sigma_{s}\hat{\pi}_{s}|^{2}\right)\mathrm{d}s+\sigma_{s}\hat{\pi}_{s}\mathrm{d}B_{s}$, we conclude that $\mathbb{E}_{t}\left[\exp\left(-p\zeta(R^{\hat{\pi}}_{T}-R^{\hat{\pi}}_{t})\right)\right]<\infty,\,a.s.$ for any $p>1$. Thus, $\hat{\pi}$ is an admissible strategy.
Next, we show that $\hat{\pi}$ is indeed an equilibrium strategy. By Lemma 3.4 and $Z^{1}=\tilde{Z}^{1}=0$, it holds that
@@
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}((\sigma_{s}\hat{\pi}_{s})^{2}-(\sigma_{s}\pi_{s}^{t,\varepsilon})^{2})\mathrm{d}s\right]$
@@
Since $r$, $\mu$, $\sigma$, $\hat{\pi}$, and $Y$ are all bounded, it suffices to show that
@@
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\hat{R}_{t},\pi^{t,\varepsilon})-J(t,\hat{R}_{t},\hat{\pi})\right)\leq\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left(-\frac{1}{2}\left(e^{-\zeta Y_{s}}+\gamma\right)(\sigma_{s}\eta)^{2}\right)\mathrm{d}s\right]\leq 0.$
Therefore, $\hat{\pi}$ is a time-consistent equilibrium strategy. ∎
@@ 4.2 Existence of time-consistent equilibrium under trading constraint
In this subsection, we assume that the trading strategy $\pi_{t}$ takes values in a bounded convex Borel set $A$. Consequently, $u_{t}=\sigma_{t}\pi_{t}$ is also bounded. To be precise, we introduce the following definition.
@@
As in the proof of Theorem 4.4, it is straightforward that $\Pi_{A}\subseteq\Pi_{0}$. Following yan2019time and liang2025integral, we next introduce the modified definitions of the perturbation strategy and the equilibrium strategy.
@@
for any $t\in[0,T)$ and any $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,A)$ . Here
$\hat{\pi}^{t,\varepsilon,a}=\left\{\begin{aligned} &a,\quad t\in[t,t+\varepsilon),\\ &\hat{\pi},\quad t\in[t+\varepsilon,T],\end{aligned}\right$.
denotes the modified perturbation strategy.
@@
$\langle w-u,\,z-u\rangle\leq 0.$
Choosing $z=u+h\in U$ yields
$\langle w-u,\,h\rangle\leq 0.$ (4.7)
It then follows that
$\langle w-u,\,v-u\rangle\geq|v-u|^{2}>0.$
Let $h=\lambda(v-u),\;\lambda\in(0,1]$ ,then
$\bigl|\alpha(w-u)-h\bigr|<\alpha|w-u|\iff 2\alpha\langle w-u,h\rangle>|h|^{2}\Longleftarrow 0<\lambda<2\alpha.$
Since $\alpha>0$, the above inequality holds for sufficiently small $\lambda>0$. By the convexity of $U$, $u+h\in U$, which completes the proof. ∎
@@
In this lemma, we consider a convex and closed set in $\mathbb{R}^{n}$ for any positive integer $n$, rather than in $\mathbb{R}$, thereby showing that our subsequent proof remains valid in the presence of multiple risky assets.
By an analogous argument in section 3, we have the following necessary condition for an equilibrium strategy.
@@
$\hat{\pi}=\sigma^{-1}\mathcal{P}_{\sigma A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}\right),\quad t\in[0,T).$ (4.8)
@@
$-\zeta\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\xi_{T}^{t,\varepsilon,a}-\hat{R}_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon,a}|^{2}\right]$
@@
Moreover, by the same argument in Lemma 3.5 and Lemma 3.6, there exists a sequence $\{\varepsilon_{n}^{t}\}$ and a fixed constant $C$ independent of $a$ such that for $a.e.$ $t\in[0,T)$,
@@
$=$ $e^{-\zeta Y_{t}}\left(\theta_{t}\Delta_{t}-\frac{1}{2}\Delta_{t}^{2}-\sigma_{t}\pi_{t}\Delta_{t}-\zeta(Z_{t}^{1}+\sigma_{t}\hat{\pi}_{t}+\rho Z_{t})\cdot\Delta_{t}\right),\quad a.s.,$
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\xi_{T}^{t,\varepsilon,a}-\hat{R}_{t})}(1-\lambda)d\lambda|\xi_{T}^{t,\varepsilon,a}|^{2}\right]\leq C\Delta_{t}^{2},\quad a.s.,$
@@
Suppose, to the contrary, that there exists some $a\in L_{\mathcal{F}_{t}}(\Omega,A)$ such that (4.11) fails to hold. Define a new admissible strategy $a_{\lambda}=\hat{\pi}_{t}+\lambda\sigma_{t}^{-1}\Delta\in A$ with sufficiently small constant $\lambda\in(0,\frac{\gamma}{\gamma+2C})$. On one hand, by (4.2), we have
@@
$\geq$ $\lambda\left(-\frac{e^{-\zeta Y_{t}}+\gamma}{2}\Delta_{t}^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\rho Z_{t}\right)-\gamma\rho\tilde{Z}-\gamma\sigma_{t}\hat{\pi}_{t}\right)\Delta_{t}\right)>0,$
@@
$\geq$ $\left|\frac{e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\left(Z_{t}^{1}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}_{t}^{1}+\rho\tilde{Z}\right)-\gamma\sigma_{t}\hat{\pi}_{t}}{e^{-\zeta Y_{t}}+\gamma}\right|,\quad a.s..$
Take $\alpha=\frac{(\zeta+1)e^{-\zeta Y_{t}}+\gamma}{e^{-\zeta Y_{t}}+\gamma}$ and $w=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}$, the above inequality can be converted to
$\left|\sigma_{t}a-\hat{u}_{t}-\alpha_{t}\left(w_{t}-\hat{u}_{t}\right)\right|\geq\left|\alpha_{t}\left(w_{t}-\hat{u}_{t}\right)\right|,\quad a.s..$
By Lemma 4.8(ii), if $\hat{u}_{t}\neq\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$, as $\alpha_{t}>1$, we can take $\lambda=1$ and then
@@
Thus we know that $\hat{u}\neq\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$ can not be an equilibrium strategy by letting $\sigma_{t}a=\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$. Therefore, (4.8) holds. ∎
As $R$ does not appear in the BSDE system, we also assume that $Z^{1}=\tilde{Z}^{1}=0$
@@
$r-\theta^{2}\leq-r-\theta\hat{u}+\frac{1}{2}\hat{u}^{2}\leq\frac{1}{2}(\hat{u}-\theta)^{2}$
Let $a_{i},i=1,2,3,4$ be (1,0),(1,0),(0,1)(0,1)(1,0),(-1,0),(0,1)(0,-1). Then $a_{i},i=1,2,3,4$ can positively span $\mathbb{R}^{2}$ and the generator satisfy the AB condition.
@@
$\left|\frac{\zeta}{2}Z^{2}+\zeta\rho Z\hat{u}-r-\theta\hat{u}+\frac{\zeta+1}{2}\hat{u}^{2}\right|\leq\frac{\zeta(1+\rho^{2})}{2}Z^{2}+k(t),$
@@
$|-r-\theta\hat{u}+\frac{1}{2}\hat{u}^{2}|\leq k(t)$
Then the BSDE system under the bounded equilibrium strategy admits a locally Holder and BMO solution (see Theorem 2.14 in xing_class_2018). Moreover, $Y$ and $\tilde{Y}$ are bounded due to the terminal condition $Y_{T}=0$.

### v2

@@ 3 A Necessary Condition of Time-Consistent Equilibria
Thus, (3.2) holds. ∎
@@ 4 Study of BSDE System in the Incomplete Factor Model
In this section, we focus on the factor model and on the two-dimensional quadratic BSDE system in (3.1), which is coupled through (3.2). The factor model provides an incomplete market setting with stochastic parameters in which the risk-free rate, the stock return rate, and the volatility rate can be expressed by a deterministic function of time $t$ and the factor $X_{t}$. We assume that the stochastic factor $X$ is governed by
@@
$\mu(t,x)>r(t,x)>0,\sigma(t,x)>0$ and $\sigma^{-1}(t,x)$ are bounded measurable deterministic functions with respect to $(t,x)$.
@@ 3. There exists a constant $L$ such that
Under Assumption 4.2, SDE (4.1) admits a unique strong solution. It is noted that the log-return process $\hat{R}$ does not appear in the BSDE system (3.1) (coupled through (3.2)), which motivates us to consider the solution adapted to the filtration $\mathbb{F}_{\bar{B}}$. We can therefore conjecture that $Z^{1}=\tilde{Z}^{1}=0$ and aim to find the solution to the following Markovian BSDE system:
$\left\{\begin{aligned} &\mathrm{d}X_{s}=m(s,X_{s})\mathrm{d}t+\nu(s,X_{s})\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}Y_{s}=\left(\frac{\zeta}{2}|\rho Z_{s}+\hat{u}_{s}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z_{s}|^{2}-a(s,\hat{\pi}_{s})\right)\mathrm{d}s+Z_{s}\mathrm{d}\bar{B}_{s},\\ &\mathrm{d}\tilde{Y}_{s}=-a(s,\hat{\pi}_{s})\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &X_{0}=x_{0},\quad Y_{T}=0,\quad\tilde{Y}_{T}=0,\\ &\hat{u}=\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}\rho Z-\gamma\rho\tilde{Z}}{(\zeta+1)e^{-\zeta Y}+\gamma},\hat{\pi}=\sigma^{-1}\hat{u},a(\cdot,\hat{\pi})=r+(\mu-r)\hat{\pi}-\frac{1}{2}|\sigma\hat{\pi}|^{2}.\end{aligned}\right$. (4.2)
Although simplified, (4.2) remains a two-dimensional quadratic BSDE system coupled through the derived equilibrium strategy $\hat{\pi}$, which prevents us from solving it in full generality. Several challenges associated with solving (4.2) in the general case are discussed in Appendix C.
In the next two subsections, we establish the existence of an equilibrium in two special cases:
  1. (i)
    When $\rho=0$, the equilibrium strategy no longer depends on (Z,Z~)($\tilde{Z}$), and the system (4.2) decouples into two separate BSDEs, for which the existence of solutions follows directly.
  2. (ii)
    Under a trading constraint, we introduce a modified definition of the equilibrium strategy. In this setting, we still obtain a BSDE system with the same form of (3.1), but the coupling strategy involves a projection operation (see (4.6)).
@@ 4.1 Existence of time-consistent equilibrium when $\rho=0$
In this subsection, we consider the case $\rho=0$. Then it follows that $\hat{\pi}$ is independent of $Z$ and $\tilde{Z}$, and thus the BSDE system (4.2) can be decoupled as
[…]
@@
$\left\{\begin{aligned} &\mathrm{d}\tilde{Y}_{s}=-r(s,X_{s})-\theta(s,X_{s})\hat{u}_{s}+\frac{1}{2}|\hat{u}_{s}|^{2}\mathrm{d}s+\tilde{Z}_{s}\mathrm{d}\bar{B}_{s},\\ &\tilde{Y}_{T}=0.\end{aligned}\right$. (4.4)
It is straightforward to see that one can first solve BSDE (4.3) and then substitute this solution into BSDE (4.4) to obtain the solution for the entire system. Regarding the existence of the solution to BSDE (4.3), we verify Conditions (2A3), (3A2), (2A1), and (4A2) and apply Theorem 4.3 in FAN20161511, which leads to the following existence and uniqueness result.
@@
Under Assumptions 4.1 and 4.2, the BSDE (4.3) has a unique solution $(Y,Z)\in\mathcal{S}^{\infty}\times L^{2}_{\mathbb{F}}(0,T;\mathbb{R})$.
@@
In the following, we verify that $\hat{\pi}$ is an equilibrium strategy by means of BSDEs (4.4) and (4.3), along with Lemma 3.4. This result is stated in the theorem below.
@@
Under Assumptions 4.1 and 4.2, when $\rho=0$, the strategy $\hat{\pi}$ defined in (4.3) is an equilibrium strategy.
@@
For notational simplicity, we adopt the same notation as in Remark 2.10 for any fixed $t\in[0,T)$, $\eta\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,\mathbb{R}^{d})$ and $\varepsilon\in(0,T-t)$.
We first verify that $\hat{\pi}\in\Pi$. Because $\hat{\pi}$ is bounded, it follows that $\hat{\pi}\in H_{\textup{BMO}}$. It is straightforward to show that
$\mathbb{E}\left[\exp\left(c\left|\int_{t}^{T}\sigma_{s}\hat{\pi}_{s}\mathrm{d}B_{s}\right|\right)\right]\leq 2\exp\left(\frac{c^{2}}{2}\|\sigma\hat{\pi}\|_{\infty}^{2}(T-t)\right),\quad\forall c>0.$
Combining this with $\hat{R}_{T}-\hat{R}_{t}=\int_{t}^{T}\left(r_{s}+(\mu_{s}-r_{s})\hat{\pi}_{s}-\frac{1}{2}|\sigma_{s}\hat{\pi}_{s}|^{2}\right)\mathrm{d}s+\sigma_{s}\hat{\pi}_{s}\mathrm{d}B_{s}$, we conclude that $\mathbb{E}_{t}\left[\exp\left(-p\zeta(\hat{R}_{T}-\hat{R}_{t})\right)\right]<\infty,\,a.s.$ for any $p>1$. Thus, $\hat{\pi}\in\Pi$.
Next, we show that $\hat{\pi}$ is indeed an equilibrium strategy. By Corollary 3.5, we obtain
@@
$=$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}(a(s,\hat{\pi}^{t,\varepsilon}_{s})-a(s,\hat{\pi}_{s})-\zeta\sigma_{s}\hat{\pi}_{s}\cdot\sigma_{s}\eta)\mathrm{d}s\right]$
$-\zeta\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\hat{\xi}_{T}^{t,\varepsilon}-\hat{R}_{t})}(1-\lambda)d\lambda|\hat{\xi}_{T}^{t,\varepsilon}|^{2}\right]$
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}(|\sigma_{s}\hat{\pi}_{s}|^{2}-|\sigma_{s}\hat{\pi}_{s}^{t,\varepsilon}|^{2})\mathrm{d}s\right]$
$\leq$ $\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}e^{-\zeta(\hat{R}_{s}+Y_{s}-\hat{R}_{t})}(a(s,\hat{\pi}^{t,\varepsilon}_{s})-a(s,\hat{\pi}_{s})-\zeta\sigma_{s}\hat{\pi}_{s}\cdot\sigma_{s}\eta)\mathrm{d}s\right]$
$+\frac{\gamma}{2}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}(|\sigma_{s}\hat{\pi}_{s}|^{2}-|\sigma_{s}\hat{\pi}_{s}^{t,\varepsilon}|^{2})\mathrm{d}s\right]$
@@
Note that $r$, $\mu$, $\sigma$, $\eta$, $\hat{\pi}$, and $Y$ are all bounded, it suffices to show that
@@
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\hat{\pi}^{t,\varepsilon})-J(t,\hat{\pi})\right)\leq\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\mathbb{E}_{t}\left[\int_{t}^{t+\varepsilon}\left(-\frac{1}{2}\left(e^{-\zeta Y_{s}}+\gamma\right)(\sigma_{s}\eta)^{2}\right)\mathrm{d}s\right]\leq 0.$
This verifies that $\hat{\pi}$ is a time-consistent equilibrium strategy. ∎
@@ 4.2 Existence of time-consistent equilibrium under trading constraint
In this subsection, we assume that the trading strategy $\pi$ takes values in a bounded, closed and convex set $A\subset\mathbb{R}$. Consequently, $u=\sigma\pi$ is also bounded. To be precise, we introduce the following definition.
@@
As in the proof of Theorem 4.4, it is straightforward that $\Pi_{A}\subseteq\Pi$. Following yan2019time and liang2025integral, we next introduce the modified definitions of the perturbed strategy and the equilibrium strategy.
@@
for any $t\in[0,T)$ and any $a\in L_{\mathcal{F}_{t}}^{\infty}(\Omega,A)$, where
$\hat{\pi}^{t,\varepsilon,a}_{s}:=\left\{\begin{aligned} &a,\quad s\in[t,t+\varepsilon),\\ &\hat{\pi}_{s},\quad s\in[0,t)\cup[t+\varepsilon,T]\end{aligned}\right$.
denotes the modified perturbed strategy.
@@
See Appendix D.8. ∎
@@
In Lemma 4.8, we consider a convex and closed set in $\mathbb{R}^{n}$ for any positive integer $n$, rather than in $\mathbb{R}$, thereby showing that our subsequent proof remains valid in the presence of multiple risky assets.
In the following discussion in this subsection, we use the similar notations as in Remark 2.10. For example, we denote the perturbed strategies $\hat{\pi}^{t,\varepsilon,a}$ and $\pi^{t,\varepsilon,a}$ simply by $\hat{\pi}^{t,\varepsilon}$ and $\pi^{t,\varepsilon}$, respectively. By an analogous argument in Section 3, we have the following necessary condition for an equilibrium strategy.
@@
$\hat{\pi}=\sigma^{-1}\mathcal{P}_{\sigma A}\left(\frac{e^{-\zeta Y}\theta-\zeta e^{-\zeta Y}(Z^{1}+\rho Z)-\gamma(\tilde{Z}^{1}+\rho\tilde{Z})}{(\zeta+1)e^{-\zeta Y}+\gamma}\right).$ (4.6)
@@
$-\zeta\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\hat{\xi}_{T}^{t,\varepsilon}-\hat{R}_{t})}(1-\lambda)d\lambda|\hat{\xi}_{T}^{t,\varepsilon}|^{2}\right]$
@@
Moreover, by the same argument in Lemma 3.6 and Lemma 3.7, there exist a sequence $\{\varepsilon_{n}^{t}\}$ and a fixed constant $C$ independent of $a$ such that for $a.e.$ $t\in[0,T)$,
@@
$=$ $e^{-\zeta Y_{t}}\left(\theta_{t}\Delta_{t}-\frac{1}{2}\Delta_{t}^{2}-\sigma_{t}\hat{\pi}_{t}\Delta_{t}-\zeta(Z_{t}^{1}+\sigma_{t}\hat{\pi}_{t}+\rho Z_{t})\cdot\Delta_{t}\right),\quad a.s.,$
$\lim\limits_{n\rightarrow\infty}\frac{1}{\varepsilon_{n}^{t}}\mathbb{E}_{t}\left[\int_{0}^{1}e^{-\zeta(\hat{R}_{T}+\lambda\hat{\xi}_{T}^{t,\varepsilon}-\hat{R}_{t})}(1-\lambda)d\lambda|\hat{\xi}_{T}^{t,\varepsilon}|^{2}\right]\leq C\Delta_{t}^{2},\quad a.s.,$
@@
Suppose, to the contrary, that there exists some $a\in L^{\infty}_{\mathcal{F}_{t}}(\Omega,A)$ such that (4.9) fails to hold. Define a new admissible perturbation $a_{\lambda}=\hat{\pi}_{t}+\lambda\sigma_{t}^{-1}\Delta\in L_{\mathcal{F}_{t}}(\Omega,A)$ with sufficiently small constant $\lambda\in(0,\frac{\gamma}{\gamma+2C})$. On one hand, by (4.2), we have
@@
$\geq$ $\lambda\left(-\frac{e^{-\zeta Y_{t}}+\gamma}{2}\Delta_{t}^{2}+\left(e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\rho Z_{t}\right)-\gamma\rho\tilde{Z}-\gamma\sigma_{t}\hat{\pi}_{t}\right)\Delta_{t}\right)>0$
@@
$\geq$ $\left|\frac{e^{-\zeta Y_{t}}\left(\theta_{t}-(\zeta+1)\sigma_{t}\hat{\pi}_{t}-\zeta\left(Z_{t}^{1}+\rho Z_{t}\right)\right)-\gamma\left(\tilde{Z}_{t}^{1}+\rho\tilde{Z}\right)-\gamma\sigma_{t}\hat{\pi}_{t}}{e^{-\zeta Y_{t}}+\gamma}\right|\quad a.s..$
Take $\alpha_{t}=\frac{(\zeta+1)e^{-\zeta Y_{t}}+\gamma}{e^{-\zeta Y_{t}}+\gamma}$ and $w_{t}=\frac{e^{-\zeta Y_{t}}\theta-\zeta e^{-\zeta Y}(Z^{1}_{t}+\rho Z_{t})-\gamma(\tilde{Z}^{1}_{t}+\rho\tilde{Z}_{t})}{(\zeta+1)e^{-\zeta Y_{t}}+\gamma}$, then the above inequality can be converted to
$\left|\sigma_{t}a-\hat{u}_{t}-\alpha_{t}\left(w_{t}-\hat{u}_{t}\right)\right|\geq\left|\alpha_{t}\left(w_{t}-\hat{u}_{t}\right)\right|,\quad a.s..$ (4.10)
Using Lemma 4.8(ii), if $\hat{u}_{t}\neq\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$, then, as $\alpha_{t}>1$, taking $\lambda=1$ yields
@@
Letting $\sigma_{t}a=\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$, we get a contradiction with (4.10). Therefore, $\hat{u}_{t}=\mathcal{P}_{\sigma_{t}A}\left(w_{t}\right)$ a.s. and hence (4.6) holds. ∎
As $R$ does not appear in the BSDE system, we can, as in the case $\rho=0$, set $Z^{1}=\tilde{Z}^{1}=0$ and derive the following BSDEs:
@@
$r-\theta^{2}\leq-r-\theta\hat{u}+\frac{1}{2}|\hat{u}|^{2}\leq\frac{1}{2}(\hat{u}-\theta)^{2}.$
Let $a_{i},i=1,2,3,4$ be (1,0),(1,0),(0,1),(0,1)(1,0),(-1,0),(0,1),(0,-1). Then $a_{i},i=1,2,3,4$ can positively span $\mathbb{R}^{2}$ and the generator satisfy the AB condition.
@@
$\left|\frac{\zeta}{2}Z^{2}+\zeta\rho Z\hat{u}-r-\theta\hat{u}+\frac{\zeta+1}{2}|\hat{u}|^{2}\right|\leq\frac{\zeta(1+\rho^{2})}{2}Z^{2}+k(t),$
@@
$|-r-\theta\hat{u}+\frac{1}{2}|\hat{u}|^{2}|\leq k(t).$
[…]